Accelerator Update for May 25, 2009 - May 27, 2009 |
Monday May 25 The day shift began with Operations monitoring the Tevatron (TeV) store 7080 (90.43E30) and the Recycler stash (304.3E10), with the Antiproton Source stacking (12.5E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the Meson MTest, NuMI, and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams. Operations terminated store 7080 at 11:45 AM. Operations put the TeV into shot setup and established store 7081 at 1:15 PM with an average initial luminosity of 308.8E30. Operations disabled the MTest beam at 6 PM. Tuesday May 26 The midnight shift began with Operations monitoring store 7081 (96.46E30) and the stash (297.2E10), with Pbar stacking (15.7E10), and with NuMI and MiniBooNE taking beam. Operations established beam to MTest at 4 AM. Operations terminated store 7081 at 5:15 AM. Operations established store 7082 at 7 AM with an initial luminosity of 310.07E30. Operations disabled the MTest beam at 6 PM. At 10:15 PM, a Booster RF station caused about 30 minutes of downtime. Operations terminated store 7082 at 11:30 PM and then put the TeV into HEP shot setup t Wednesday May 27 The midnight shift began with the TeV in shot setup, with Operations monitoring the stash (384.7E10), with Pbar stacking (33.5E10), and with NuMI and MiniBooNE taking beam. The HEP shot setup ended with a TeV quench at D0. A D0 quench protection equipment problem was investigated during the rest of midnight shift. The Plans for Wednesday The plans for today are to complete the TeV checkout and then resume stacking, stashing, and storing. |
|
More Information Comments and Suggestions |
| last modified 05/28/2009 email Fermilab |
| Security, Privacy, Legal | |